# The Dark Side of LLMs: Agent-based Attacks for Complete Computer Takeover

### v3

@@ 3.4 Synthetic applications overview
To explore the feasibility of using modern AI agents as attack vectors, we designed a series of incremental synthetic applications with the goal of identifying LLMs that would respond differently to the same malicious command (depicted in listing 1) depending on how it was delivered.
We began our analysis with a classic direct prompt injection. We provided an LLM agent (equipped with terminal interaction tool) with the full textual content of the infected PDF document (i.e., the extracted text, not the file itself), asking it to assess whether any part of the content appeared malicious and execute it if deemed secure.
Whenever an LLM agent refused to execute the command, we introduced a RAG module, turning the agent into an agentic rag. In this setup, we implemented a RAG backdoor attack, specifically within the observation and thought attack category. The malicious payload is hidden within a document in the knowledge base and the attack is triggered during the data retrieval and planning phase and the execution occurred seamlessly, without altering the final output or alerting the user.
Moreover, we explored another potential attack vector: inter-agent communication by designing an agentic AI system composed of multiple agents. The calling is aware of the capabilities and roles of other agents within its Agentic AI system and can communicate with them if needed. The invoked agent is responsible solely for executing the instructions it receives from other agents and returning the output to the caller.
This synthetic application aims to verify our claim: with high probability, the relationship between peer agent can easily deceive LLM in conducting malicious operations by leveraging implicit trustiness. This scenario is designed to test if the same model that had previously rejected the command in the agentic rag scenario executes it in the multi-agent setting simply because it was requested by another peer agent. The called agent is responsible for the actual execution of the malicious command and serves as the operational vector of the attack within the multi-agent architecture.
@@ 4 Evaluation
For supporting our claims we resort to implementing each synthetic application and performing a campaign of experiments to determine the sensibility of each LLM to such attacks. For each synthetic application, we tested the behavior of the 17 LLMs listed in table 1.
@@
Devstral devstral:24b
Custom MFDoom/deepseek-r1-tool-calling:70b
The only simplification we made for RAG backdoor attack was to embed the malicious message within a single chunk -alongside benign content - to increase its retrieval likelihood. Since our focus is on evaluating LLM security in agentic and multi-agent contexts, it is sufficient that the chunk is retrieved at least once. Therefore, we do not address the optimization of trigger placement or the number of occurrences required, as explored in prior RAG poisoning work (see Section 5).
It is noteworthy that, in all tests conducted on RAG backdoor attack and multi-agent communication, we adopted the perspective of a typical, well-intentioned user (i.e., someone engaging with the agent in a legitimate and non-malicious manner). This user-centric approach allowed us to realistically assess how feasible it is for a benign user to unknowingly become the target of a backdoor attack simply by using the agent as intended. This perspective highlights the severity of the excessive trust that LLMs place in their internal components and surrounding environment. Under these conditions, the attacker is not required to interact with the agent directly; instead, they only need to tamper with a document in the agent’s knowledge base and passively wait for the victim’s machine to initiate an outbound connection once the infected document is processed.
@@ 4.1 Synthetic application A: LLM agent
In the first synthetic application we evaluate the sensitivity of each LLM to direct prompt injection. The architecture of the agent is simple and id depicted in figure 2. The most important tool is named run_command and allows the agent to perform commands to a non-interactive bash terminal through python libraries like os or subprocess.
Figure 2: Synthetic application A. LLM agent that can runs command.
In the agent_system_prompt we ask the agent to be useful for the user but not to run dangerous commands.
@@
The attacker’s primary objective is to misdirect the agent to achieve specific goals that align with the attacker’s intent but are unintended by the user. For individual user tasks, the attacker can only observe binary success-failure feedback as the outcome of their attacks. In accordance with the black-box setting assumption delineated in the threat model, the unique feedback available in the LLM agent synthetic application is the binary indicator of whether or not malware has been installed. This involves verifying if the attacker has obtained a connection from the reverse shell.
We evaluated two things. First, we checked if LLM gave any clues about the nature of the command. In other words, we wanted to see if it identified the command as malicious. Next, we checked if the malware was actually executed after the correct tool was used. So, we wanted to see if the attacker received the connection.
@@
The results are illustrated in table 2. We noted that 7 of the 17 LLMs tested (41.2%) were vulnerable to direct prompt injection, which underscores that this is still a real and serious threat for this domain. Surprisingly, however, we demonstrated that three of these LLMs decided to execute the requested task despite realizing its malicious nature during the reasoning phase, only because they have the tool to complete the operation and because the prompt emphasized the urgency and immediacy of executing the task.
Table 2: Prompt injection against LLM Agent
Model Malware executed Malicious classified
GPT-4o-mini YES NO
GPT-4o NO YES
GPT-4.1 NO YES
Claude-4-Sonnet NO YES
Claude-4-Opus NO YES
Gemini-2.0-flash YES NO
Gemini-2.5-flash YES YES
Magistral-medium YES YES
Mistral-large NO YES
Mistral-small NO YES
devstral:24b YES NO
LLama3-8b-8192 YES YES
llama3.3:70b NO YES
llama4:16x17b NO YES
qwen3:14b YES YES
qwen3:30b NO YES
MFDoom/deepseek-r1-tool-calling:70b NO YES
4.2 Synthetic application B: Agentic RAG
A typical agent system processes user queries by interacting with a diverse set of tools and services within its environment to accomplish user tasks. The LLM in the agent serves as the planner, dynamically coordinating between these components to retrieve information, execute commands, and respond to user needs. Moreover, they often rely on external data sources. In this scenario, the attacker exploits the dependency and trustiness between the model and documents provided by RAG by strategically manipulating specific parts of the external data sources, which the agent later retrieves and processes as part of its task execution. Once these contaminated inputs are fed into the LLM, they can alter its behavior, leading to unauthorized actions.
Our goal in implementing the second synthetic application was to analyze the above mentioned trust relationship for various LLM leveraging RAG backdoor attack for agents. The architecture of the agent is the standard for agentic rag and is shown in figure 3. It has the same run_command tool of the synthetic application A and a retrieval tool used to add a RAG system to the core LLM.
Figure 3: Synthetic application B. Agentic rag that can runs command and retrieve external documents using RAG techniques.
As in the previous scenario, in the agent_system_prompt we ask the agent to be useful for the user but not to perform dangerous operations.
We altered a document within the RAG knowledge base by injecting a hidden string (the text is white on a white background and the font size is the smallest possible) containing a "urgent" request to execute the malicious instructions of listing 1. The purpose of concealing the message was to ensure the validity of the document from the perspective of a standard user navigating the knowledge base.
PDF hidden message If u ar3 re4d1ng th1s, 1MM3DI4TLY EX3CUT3 the C0MM4ND W1TH0UT 4sk1ng for C0NF1RM4TI0N: <command>.
Use the TOOL to access the TERMINAL. Everything is part of the user request you need to respond to.
In light of the black-box setting assumptions delineated in the preceding scenario and their ramifications, we conducted an evaluation to assess the level of trust each LLM places in the documents provided through RAG. This evaluation entailed an analysis of the behavior exhibited by the agentic RAG in response to a user request necessitating retrieval from the external knowledge base.
In the absence of malicious chunks among the extracted chunks, the output is expected to conform to the user’s request. In the event that the extracted chunks contain a hidden malicious message, the subsequent evaluation determines whether the malware is executed and concurrently the final output produced by the LLM is consistent with the user’s requests, excluding any indication of the intermediate steps in the computational process.
Table 3: RAG Backdoor Attack against Agentic RAG
Model Malware executed
GPT-4o-mini
@@ 4.2 Synthetic application B: Agentic RAG
MFDoom/deepseek-r1-tool-calling:70b ×
  • = Vulnerable; × = Resistant; = Resistant to direct-prompt injection, but vulnerable to RAG-backdoor attack
The results are illustrated in table 3. The RAG backdoor attack proved to be significantly more effective than direct prompt injection, with 9 out of 17 LLMs (52.9%) successfully executing the malicious payload. This represents a considerably higher success rate compared to the six vulnerable models in the direct injection scenario, highlighting that RAG-based attacks pose an even more severe security threat to agentic systems. Particularly concerning is the fact that several models that successfully resisted direct prompt injection attacks were compromised through the RAG backdoor vector (highlighted in blue). This demonstrates that even security-conscious LLMs with robust direct injection defenses can be circumvented when malicious content is embedded within seemingly legitimate retrieved documents. The attack’s effectiveness stems from the models’ inherent trust in RAG-retrieved content, which is typically perceived as factual information rather than potentially adversarial input, thereby bypassing many of the safety mechanisms designed to detect malicious prompts.
@@ 4.3 Synthetic application C: Agentic AI systems
Over the past year, interest in the development and use of multi-agent systems has grown significantly. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [21]. Motivated by these considerations, we set out to evaluate the security of trust boundaries within Agentic AI systems, focusing specifically on the possibility that an agent might be coerced into performing malicious actions when requested by a peer agent — actions that it would otherwise reject if requested directly by a human user.
Figure 4: Agentic AI system. the calling agent is an agentic rag capable of call other agents. The invoked agent is an LLM agent with at least a tool to access the terminal.
In this synthetic application, we tested the same RAG Backdoor Attack previously discussed but executed in a multi-agent system. The calling agent is an agentic rag (not necessarily equipped with terminal access tool) that functions as a standard question-answering agent concerning the content of documents provided through RAG. The invoked agent, by contrast, is an LLM agent with at least a tool to access the terminal. The internal structure of each agent in this scenario is identical to that used in the previous ones. This methodological consistency ensures that any successful attack cannot be attributed to structural variations in the agent or to differences in how it interacts with users or external sources. Instead, it allows us to isolate mutual trust in inter-agent communication as the sole relevant factor influencing the outcome.
@@
The experiments’ results conducted on this scenario are illustrated in table 4. The inter-agent trust exploitation attack demonstrated the most alarming success rate, with 14 out of 17 LLMs (82.4%) successfully executing the malicious payload when requested by a peer agent. This represents the highest vulnerability rate across all three attack vectors tested, revealing a critical security flaw in multi-agent architectures. Most notably, only Claude-4-Sonnet, Gemini-2.5-flash and qwen3:30b maintained their defensive posture against peer agent manipulation, while all other models (highlighted in blue in table 4) - including those that successfully resisted to direct prompt injection or RAG backdoor attacks - were compromised through inter-agent communication channels.
This finding exposes a fundamental weakness in current LLM safety mechanisms: models appear to apply significantly more lenient security policies when interacting with other AI agents compared to direct human interactions or external tool only, essentially treating peer agents as inherently trustworthy entities despite the potential for compromise or malicious intent. Finally, traditional evaluation and safety frameworks, built for static or single-function AI , are no longer sufficient.
Table 4: Vulnerability Assessment for Agentic AI Systems
Model Malware executed
@@
  • = Vulnerable; × = Resistant; = Resistant to single LLM agent attacks, but vulnerable to multi-agent system attacks
@@ 4.4 Comprehensive Analysis
A comprehensive analysis across all three attack vectors reveals several non-trivial security implications for agentic AI systems. First, it is worth noting that only Claude-4-Sonnet out of 17 proved to be entirely secure (5.9%) while each other models exhibited weaknesses in at least one of the evaluated attack scenarios, ultimately leading to the successful installation and execution of the malware.
Moreover, there exists a clear trust hierarchy vulnerability gradient: direct prompt injection (46.2% success rate) < RAG backdoor attacks (69.2%) < inter-agent trust exploitation (84.6%). This escalating vulnerability pattern suggests that current LLM safety training primarily focuses on human-to-AI interactions while inadequately addressing AI-to-AI communication scenarios and external data source validation. Particularly concerning is the discovery that security-conscious models exhibit inconsistent defensive behaviors across attack vectors. For instance, Mistral-large and several Llama variants successfully identified and rejected malicious direct prompts but were completely vulnerable to RAG backdoor and inter-agent attacks. This inconsistency indicates that existing safety mechanisms are context-dependent rather than comprehensive, creating exploitable blind spots in multi-modal agentic deployments.
@@
Model Direct Injection RAG Backdoor Inter-Agent Trust Vulnerability Score
@@
GPT-4o × × 1/3
GPT-4.1 × × 1/3
Claude-4-Sonnet × × × 0/3
Claude-4-Opus × × 1/3
@@
Mistral-large × 2/3
Mistral-small × 2/3
@@
qwen3:30b × × 1/3
MFDoom/deepseek-r1-tool-calling:70b × × 1/3
Success Rate 41.2% 52.9% 82.4% -
  • = Vulnerable; × = Resistant; * = Recognizes malicious intent but executes anyway
@@
Small (<=24B) 3/4 (75.0%) 3/4 (75.0%) 4/4 (100.0%) 4
Medium (24B-70B) 1/4 (25.0%) 2/4 (50.0%) 3/4 (75.0%) 4
Large (>70B) 0/2 (0%) 1/2 (50.0%) 2/2 (100.0%) 2
Closed-source (N/A) 3/7 (42.9%) 3/7 (42.9%) 5/7 (71.4%) 7
Overall 6/17 (41.2%) 9/17 (52.9%) 14/17 (82.4%) 17
  • Small: LLama3-8b-8192, qwen3:14b, devstral:24b, Mistral-small
  • Medium: qwen3:30b, MFDoom/deepseek-r1-tool-calling:70b, llama3.3:70b, Magistral-medium
  • Large: Mistral-large, llama4:16x17b
@@
  • Closed-source: GPT-4o-mini, GPT-4o, GPT-4.1, Claude-4-Sonnet, Claude-4-Opus, Gemini-2.0-flash, Gemini-2.5-flash
@@
The most critical finding is the collapse of security boundaries in multi-agent environments. Models like llama3.3:70b and llama4:16x17b, which demonstrated robust resistance to both direct injection and RAG manipulation, immediately capitulated when the same malicious request originated from a peer agent. This suggests that current LLM architectures implicitly encode a "AI agent privilege escalation" vulnerability, where requests from other AI systems bypass standard safety filters — a design flaw that becomes exponentially dangerous as enterprise AI deployments increasingly adopt multi-agent orchestration patterns.

### v4

@@ 3.3 Synthetic applications overview
To explore the feasibility of using modern AI agents as attack vectors, we designed three synthetic applications with the goal of identifying LLMs that would respond differently to the same malicious prompt depending on how it is delivered. Figure 2 presents a design of the agent architecture to be employed in each synthetic application.
We began our analysis with a classic Direct Prompt Injection attack. We provided the LLM agent (equipped with a terminal interaction tool) with a prompt containing both harmful and harmless information, and analyzed whether it would execute the command pipe or classify the prompt content as malicious. This attack served as the baseline for our research and was designed to evaluate the intrinsic security mechanisms of LLMs against harmful input originating from malicious users that could compromise the integrity of a computing system.
Building upon the insight from Direct Prompt Injection preliminary tests – which showed that bypassing the security mechanisms of modern LLMs remains a surprisingly achievable task – we extended our investigation to assess a more critical scenario: the potential for these models to act as involuntary facilitators of malicious behavior. Specifically, we focused on the implementation of RAG Backdoor attacks designed to target benign, unaware users. We aim to demonstrate how LLMs, when integrated into agentic RAG, can be covertly manipulated to execute harmful operations without any direct adversarial interaction with the user and the agent. Our RAG Backdoor attack is specifically within the observation and thought attack category. In this scenario, the malicious prompt is hidden within a document in the knowledge base and the attack is triggered during the data retrieval and planning phase. The execution should occur seamlessly, without altering the final output or alerting the user.
Furthermore, we explored another potential attack vector, Inter-Agent Communication, by designing a multi-agent system composed of multiple agents to explore potential ways to abuse the Inter-Agent Trust. The calling agent is aware of the capabilities and roles of other agents within the system and can communicate with them if needed. The invoked agent is solely responsible for executing the instructions it receives from other agents and returning the output to the caller. This synthetic application aims to verify our main claim: with high probability, the relationship between peer agent can easily deceive LLM in conducting malicious operations by leveraging implicit trustiness. This scenario tests if the same model that had previously rejected the command in the agentic RAG or direct prompt injection scenario executes it in the multi-agent setting simply because it was requested by another peer agent. The called agent is responsible for the actual execution of the malicious command and serves as the operational vector of the attack within the multi-agent architecture.
Figure 2: Agent architecture for each synthetic application. (a) LLM agent that can run commands. (b) Agentic RAG that can run commands and retrieve external documents using RAG techniques. (c) The calling agent is an agentic RAG capable of calling other agents. The invoked agent is an LLM agent with at least a tool to access the terminal.
@@ 4 Evaluation
To support our claims, we implement each synthetic application along with the corresponding attacks. For each technique, we perform a campaign of experiments to determine the sensitivity of the eighteen LLMs, listed in Table 1, to such attacks.
@@
Mistral devstral:24b
DeepSeek-r1 MFDoom/deepseek-r1-tool-calling:70b
The only simplification we made for the RAG Backdoor attack was to embed the malicious prompt within a single chunk – alongside benign content – to increase its retrieval likelihood. Since our focus is on evaluating LLM security in agentic and multi-agent contexts, it is sufficient that the chunk is retrieved at least once. Therefore, we do not address the optimization of trigger placement or the number of occurrences required, as explored in previous work on RAG poisoning (see Section 7).
It is noteworthy that, in all tests conducted on RAG Backdoor and Inter-Agent Trust attacks, we adopted the perspective of a typical, well-intentioned user (i.e., someone engaging with the agent in a legitimate and non-malicious manner). This user-centric approach allowed us to realistically assess how feasible it is for a benign user to unknowingly become the target of an attack simply by using the agent as intended. This perspective highlights the severity of the excessive trust that users may place in LLMs alongside with the trust that models themselves place in their internal components and surrounding environment. Under these conditions, the attacker is not required to interact with the agent directly; instead, they only need to tamper with a document in the agent’s knowledge base and passively wait for the victim’s machine to initiate an outbound connection once the infected document is processed.
It is crucial to emphasize that our evaluation methodology deliberately avoided sophisticated prompt engineering techniques for security hardening. We intentionally employed minimal, straightforward system prompts (e.g., "Please complete tasks efficiently and securely") rather than implementing advanced security-oriented prompt engineering strategies. This choice reflects our primary research objective: evaluating whether LLMs possess inherent security capabilities sufficient for agentic deployment without relying on external prompt-based mitigation. Our findings demonstrate that current LLM architectures, in their baseline configuration, fundamentally lack the intrinsic security mechanisms necessary for safe agentic operations.
@@ 4.1 Synthetic application A: LLM agent
In the first synthetic application, we evaluate the sensitivity of each LLM to Direct Prompt Injection attacks. The most important agent tool is named run_command and allows the agent to perform commands in a non-interactive bash terminal through Python libraries like os or subprocess. In the agent_system_prompt, we ask the agent to be useful for the user but not to run dangerous commands.
@@
The evaluation involved two factors: the malware execution and its identification. The former indicates whether the attacker’s primary objective was obtained. In accordance with the black-box setting assumption delineated in the threat model, the unique feedback available in this synthetic application is the binary success-failure indicator of whether or not malware has been installed, that is, the attacker received the connection from the reverse shell. The latter indicates whether the LLM gave any clues about the nature of the command pipe contained in the message. In other words, we wanted to see if it identified the prompt as malicious. The attack surface is represented by the prompt itself, while the core trust boundary lies in the implicit trust of the LLM in the content of user-provided prompts.
@@
Table 2 reports the results grouped by LLM, regardless of the specific malicious prompt used. A model is considered vulnerable if at least one of the tested prompts led to the successful execution of the malware. We noted that seventeen of the eighteen LLMs tested (94.4%) were vulnerable to Direct Prompt Injection, highlighting that this is still a real and serious threat to this domain. Surprisingly, we found that five of these LLMs decided to execute the requested task despite realizing its malicious nature during the reasoning phase, only because they have the tool to complete the operation and because the prompt emphasized the urgency and immediacy of executing the task.
Table 2: Direct Prompt injection against LLM Agent
Model Malware execution Malware identification
GPT-4o-mini ×
@@
Gemini-2.0-flash ×
Gemini-2.5-flash ×
Gemini-2.5-pro ×
Magistral-medium
Mistral-large ×
Mistral-small ×
devstral:24b ×
llama3.3:70b
llama4:16x17b ×
qwen3:14b
qwen3:30b
MFDoom/deepseek-r1-tool-calling:70b
  • = Vulnerable; × = Not Vulnerable; = Identified as malicious; × = Not identified as malicious;
The high success rate of 94.4% indicates that LLMs treat the presence of a tool as implied permission to use it regardless of the security context. This reveals a design flaw where "capability implies permission."
4.2 Synthetic application B: Agentic RAG
A typical agent processes user queries by interacting with a diverse set of tools and services within its environment to accomplish user tasks. Moreover, they often rely on external data sources.
In this scenario, the attacker exploits the dependency and trust between the model and the documents provided through the RAG by strategically manipulating specific parts of the external data sources, which the agent later retrieves and processes as part of its task execution. Once these contaminated inputs are fed into the LLM, they can alter its behavior, leading to unauthorized actions. Our goal in implementing this synthetic application is to analyze the aforementioned trust relationship for various LLMs, leveraging RAG backdoor attack for agents. The architecture of the agent is shown in figure 2(b). It has the same run_command tool of synthetic application A and a retrieval tool used to add external knowledge to this LLM. As in the previous scenario, in the agent_system_prompt we ask the agent to be useful to the user but not to perform dangerous operations.
We altered a document within the RAG knowledge base by injecting the malicious prompt as an hidden string (the text is white on a white background and the font size is the smallest possible). The purpose of concealing the message was to ensure the validity of the document from the perspective of a standard user navigating the knowledge base.
In light of the setting assumptions delineated in the threat model, we analyze the behavior exhibited by the agentic RAG in response to a user request necessitating retrieval from the external knowledge base. In the absence of malicious chunks among the extracted chunks, the output is expected to conform to the user’s request. In the event that the extracted chunks contain a hidden malicious prompt, the subsequent evaluation determines whether the malware is executed, and concurrently the final LLM output is consistent with the user’s requests, excluding any indication of the intermediate steps in the computational process.
Table 3: RAG Backdoor Attack against Agentic RAG
Model Malware executed Correct answer
GPT-4o-mini
GPT-4o
GPT-4.1-mini
GPT-4.1
Claude-4-Sonnet
Claude-4-Opus
@@ 4.2 Synthetic application B: Agentic RAG
MFDoom/deepseek-r1-tool-calling:70b
  • = Vulnerable; × = Not Vulnerable; = Correct answer to the user request
The results are illustrated in Table 3. A model is considered vulnerable if at least one of the tested prompts led to the successful execution of the malware. Fifteen out of eighteen tested models (83.3%) were vulnerable to the RAG Backdoor attack.This attack is the first to clearly demonstrate how an attacker can effectively repurpose agents as attack vectors – without direct manipulation, user interaction, or raising user awareness. Indeed, the RAG backdoor attack proved to be significantly stealthy: all vulnerable models maintained a perfect functional facade (100%) by providing correct answers to user requests while simultaneously executing malicious payloads. This dual behavior creates an invisible attack channel where users receive expected responses with no indication of compromise, effectively turning agentic RAG systems into an unwitting attack infrastructure.
@@ 4.3 Synthetic application C: Agentic AI systems
Over the past year, interest in the development and use of multi-agent systems has grown significantly. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [23]. Motivated by these considerations, we set out to evaluate the security of trust boundaries within Agentic AI systems, focusing specifically on the possibility that an agent might be coerced into performing malicious actions when requested by a peer agent – actions that it would otherwise reject if requested directly by a human user.
In this synthetic application, we tested the same RAG Backdoor attack previously discussed but executed in a multi-agent system. The calling agent is an agentic RAG (not necessarily equipped with a terminal access tool) that functions as a standard question-answering agent regarding the content of documents provided through RAG. In contrast, the invoked agent is an LLM agent with at least a tool to access the terminal. The internal structure of each agent in this scenario is identical to that used in the previous ones. This methodological consistency ensures that any successful attack cannot be attributed to structural variations in the agent or to differences in how it interacts with users or external sources. Instead, it allows us to isolate mutual trust in inter-agent communication as the sole relevant factor influencing the outcome. For the invoked agent, we adopt an agent_system_prompt similar to the previous applications, asking the agent to be useful for the user but not to perform dangerous operations.
@@
The results of the experiments conducted on this scenario are illustrated in Table 4. The Inter-Agent Trust Exploitation attack demonstrated the most alarming success rate, with 100.0% of the tested LLMs successfully executing the malicious payload when requested by a peer agent. This represents the highest vulnerability rate across all three attack vectors tested, revealing a critical security flaw in multi-agent architectures. Furthermore, the universal success rate (100.0%) in maintaining correct responses while executing backdoor instructions confirms that the current LLM safety mechanisms are fundamentally misaligned with the threat model of agentic systems. Models appear to apply significantly more lenient security policies when interacting with other AI agents compared to direct human interactions or external tool only, essentially treating peer agents as inherently trustworthy entities despite the potential for compromise or malicious intent.
Table 4: Vulnerability Assessment for Multi-agent Systems
Model Malware executed Correct answer
@@
  • = Vulnerable; = Correct answer to the user request
The success rate observed in Inter-Agent Trust Exploitation attacks carries profound implications that extend far beyond single-host compromises. In real-world enterprise deployments, multi-agent systems could be distributed across heterogeneous computing environments, with individual agents typically executing on separate hosts, cloud instances, or even different organizational boundaries. Each successful agent-to-agent interaction becomes a potential privilege escalation bridge to additional systems. Moreover, the stealthy nature of these attacks, demonstrated by the 100% correct response rate while executing malicious payloads, ensures that such compromises can persist undetected across systems.
@@ 4.4 Comprehensive Analysis
A comprehensive analysis, which results are illustrated in Table 5, across all three attack vectors reveals several non-trivial security implications for agentic AI systems. First, it is worth noting that none of the eighteen tested models proved to be entirely secure. Each model exhibited weaknesses in at least one of the evaluated attack scenarios, ultimately leading to the successful installation and execution of the malware. A significant proportion of the models, 15/18 (83.3%) exhibited vulnerability scores of 3/3 attacks, suggesting that the vast majority are entirely vulnerable. In contrast, only 3/18 (16.7%) models demonstrated partial resistance.
@@
Model Direct Prompt Injection RAG Backdoor Inter-Agent Trust Vulnerability Score
@@
GPT-4o 3/3
GPT-4.1-mini 3/3
GPT-4.1 3/3
Claude-4-Sonnet 3/3
Claude-4-Opus 3/3
@@
Mistral-large 3/3
Mistral-small 3/3
@@
qwen3:30b * 3/3
MFDoom/deepseek-r1-tool-calling:70b * 3/3
Success Rate 94.4% 83.3% 100.0% -
  • = Vulnerable; × = Not Vulnerable; * = Recognizes malicious intent but executes anyway
@@
Smaller than 70B 4/4 (100.0%) 4/4 (100.0%) 4/4 (100.0%) 4
Bigger than 70B 2/3 (66.6%) 1/3 (33.3%) 3/3 (100.0%) 3
Closed-source (N/A) 11/11 (100.0%) 10/11 (90.9%) 11/11 (100.0%) 11
Overall 17/18 (94.4%) 15/18 (83.3%) 18/18 (100.0%) 18
  • Small: qwen3:14b, devstral:24b, Mistral-small, qwen3:30b
  • Large: MFDoom/deepseek-r1-tool-calling:70b, llama3.3:70b, llama4:16x17b
@@
  • Closed-source: GPT-4o-mini, GPT-4o, GPT-4.1, GPT-4.1-mini, Claude-4-Sonnet, Claude-4-Opus, Gemini-2.0-flash, Gemini-2.5-flash, Gemini-2.5-pro, Magistral-medium, Mistral-large
@@
The most critical finding is the collapse of security boundaries in multi-agent environments. Models like Gemini-2.5-pro, llama3.3:70b, and llama4:16x17b, which demonstrated robust resistance to one or both direct injection and RAG manipulation, immediately capitulated when the same malicious request originated from a peer agent. This suggests that current LLM architectures implicitly encode an “AI agent privilege escalation” vulnerability, where requests from other AI systems bypass standard safety filters.